# MCTS-SQL: An Effective Framework for Text-to-SQL with Monte Carlo Tree Search

### v1

@@ Abstract
Text-to-SQL is a fundamental and longstanding problem in the NLP area, aiming at converting natural language queries into SQL, enabling non-expert users to operate databases. Recent advances in LLM have greatly improved text-to-SQL performance. However, challenges persist, especially when dealing with complex user queries. Current approaches (e.g., COT prompting and multi-agent frameworks) rely on the ability of models to plan and generate SQL autonomously, but controlling performance remains difficult. In addition, LLMs are still prone to hallucinations. To alleviate these challenges, we designed a novel MCTS-SQL to guide SQL generation iteratively. The approach generates SQL queries through Monte Carlo Tree Search (MCTS) and a heuristic self-refinement mechanism are used to enhance accuracy and reliability. Key components include a schema selector for extracting relevant information and an MCTS-based generator for iterative query refinement. Experimental results from the SPIDER and BIRD benchmarks show that MCTS-SQL achieves state-of-the-art performance. Specifically, on the BIRD development dataset, MCTS-SQL achieves an Execution (EX) accuracy of 69.40% using GPT-4o as the base model and a significant improvement when dealing with challenging tasks, with an EX of 51.48%, which is 3.41% higher than the existing method.
Figure 1: An example of Text-to-SQL, where the user’s queries are highly complex, leading to challenges for the LLM in accurately selecting the appropriate fields and solving tasks such as table selection and joins.
@@ 1 Introduction
Text-to-SQL aims at converting natural language queries to SQL, which plays a critical role in data analytics and supports a wide range of applications (e.g., business intelligence and automated reporting). Recent advances in LLM have significantly improved the accuracy and stability of text-to-s SQL systems.
However, real-world Text-to-SQL applications face with significant challenges, mainly due to user queries’ random, often ambiguous nature (as seen in Figure 1). Text-to-SQL is challenging for three main reasons. First, natural language has ambiguous expressions and rich semantics, making it hard to map them to specific database tables and fields. Second, extracting meaningful information from natural language requires strong understanding capabilities. Lastly, generating accurate SQL queries that meet often vague requirements demands both semantic understanding and syntactic precision.
To tackle these challenges, existing algorithms have employed various strategies, typically falling into categories such as rule-based methodsKang et al. (2012); Hammami et al. (2021), sequence-to-sequence modelsZhong et al. (2017); Katsogiannis-Meimarakis and Koutrika (2021), and pre-trained language modelsKatsogiannis-Meimarakis and Koutrika (2023); Guo and Gao (2019). Early Text-to-SQL systems relied on rule-based approaches, which employed handcrafted grammar rules and heuristics. Subsequently, sequence-to-sequence models and pre-trained language models were introduced, enabling more accurate and flexible SQL generation. In the past three years, the advent of LLMs has significantly advanced Text-to-SQL, particularly in zero-shot and few-shot settings where models can generate SQL without extensive training examples.Despite these advancements, challenges remain, such as handling ambiguous queries and optimizing SQL for both performance and correctness. Most methods use COT or multi-agent frameworks, but due to the hallucination in LLMs, the ability of these self-planning strategies to solve complex queries is difficult to control.
A promising way to alleviate these limitations is to design a controllable chain of thought reasoning path, rather than allowing LLMs to generate it autonomously. Recent researchesPitanov et al. (2023); Li et al. (2023); Chen et al. (2024) have shown that Monte Carlo Tree Search(MCTS) is effective in complex decision-making and optimization tasks, making it a promising tool for Text-to-SQL. Through multiple iterations, MCTS explicitly selects the path with the highest probability to solve complex problems. In this paper, we integrate MCTS with Text-to-SQL, developing a more robust framework of tackling challenging SQL generation tasks that current LLMs struggle with. Furthermore, MCTS can effectively alleviate the mapping of ambiguous semantics in natural language to database fields through continuous attempts and feedback, thereby enhancing the alignment between the generated SQL and the user’s vague intent.
However, MCTS typically requires a large number of additional tokens, which is unnecessary for simple SQL generation tasks. Therefore, our framework is designed with fast-slow thinking modules, which comprise a direct SQL generation component and an MCTS-based Refiner. Only when the SQL generation module fails to generate a valid query, the MCTS-based Refiner is employed for iterative optimization. The system performs a self-evaluation of SQL generation errors and, based on the results, applies Monte Carlo Tree Search to choose a better path to refine the solution step by step. Additionally, we propose a selector module, which partitions large databases into smaller sub-databases, identifying the most relevant data tables and columns.
@@
  • To our best knowledge, we are the first to apply Monte Carlo Tree Search(MCTS) to the Text-to-SQL domain, capitalizing on its superior ability to optimize decision-making in complex and challenging queries.
  • We introduce MCTS-SQL, an innovative Text-to-SQL framework that integrates a dual-process thinking model, comprising a fast SQL generation module and an MCTS-based refinement process, with an auxiliary selector to optimize the overall performance.
  • Through comprehensive experiments, we highlight the complementary strengths of LLMs and MCTS, demonstrating substantial improvements in handling complex Text-to-SQL tasks.
  • In our experiments, we conducted extensive testing on the widely used Spider and BIRD datasets. The results demonstrate that MCTS-SQL, based on GPT-4, achieved a query execution accuracy of 69.40% on the holdout dev set of BIRD, establishing state-of-the-art performance, particularly with significant improvements on challenging tasks, where it reaches 51.48%, outperforming existing methods by 3.41%.
@@ 2.1 Text-to-SQL
Text-to-SQL aims to bridge natural language queries and structured database queries. Due to its complexity, numerous approaches are proposed to address its challenges. Early systems, such as LUNARKang et al. (2012) and NaLIXHammami et al. (2021), employed rule-based methods that manually crafted grammar rules and heuristics. However, these approaches are highly domain-specific, and generalization performance across different tasks or databases is difficult to guarantee.
The deep learning marked a turning point for Text-to-SQL. End-to-end models like Seq2SQLZhong et al. (2017) and SQLNetKatsogiannis-Meimarakis and Koutrika (2021) directly mapped natural language to SQL but struggled with complex queries, especially those involving nested structures or intricate reasoning. Pre-trained Language Models (PLMs), such as TaBERTKatsogiannis-Meimarakis and Koutrika (2023) and BERT-SQLGuo and Gao (2019), enhance cross-domain generalization and improve the accuracy of SQL generation. However, handling complex user queries and achieving powerful cross-domain performance remain significant challenges.
Recently, Large Language Models (LLMs) such as GPT-4Achiam et al. (2023), Palm-2Anil et al. (2023), and LLaMATouvron et al. (2023) have revolutionized Text-to-SQL tasks. These models excel in zero-shot and few-shot settings due to extensive pretraining and advanced reasoning capabilities.Lee et al. (2024); Talaei et al. (2024); Alp Caferoğlu and Ulusoy (2024) Techniques like DAIL-SQLGao et al. (2023) optimized prompt ngineering, focusing on question representation, prompt structure, and example selection to enhance SQL accuracy with minimal supervision. Frameworks like C3-SQLDong et al. (2023), DIN-SQLPourreza and Rafiei (2024), and StructGPTJiang et al. (2023) further advanced the field by addressing complex queries through database simplification, query decomposition, and structured data access. Additionally, MAC-SQLWang et al. (2024) introduced a collaborative framework integrating decomposer, auxiliary selector, and refiner modules for iterative SQL refinement.
From the analysis of existing methods, it is evident that advancing Text-to-SQL hinges on improving models’ ability to reason through complex scenarios. Current approaches often rely on basic COT strategies or multi agent framework to autonomously plan, yet they suffer from hallucination issues. To address this, our work integrates Monte Carlo Tree Search to enhance models’ comprehension of ambiguous table structures and queries, significantly improving SQL generation accuracy.
Figure 2: The MCTS-SQL framework is structured around three primary components: the Selector, responsible for extracting relevant information across diverse tables and columns, and the generator (consisting of the Direct Generator and MCTS-Refiner), a hybrid module inspired by the principles of fast and slow thinking. The Generator integrates a direct SQL generation component with an MCTS-based Refiner. In cases where the direct SQL generation fails to address the query effectively, the MCTS-based Refiner performs iterative.
@@ 2.2 Monte Carlo Tree Search
MCTS is a widely used tool for planning complex problems, and a large number of downstream experiments have demonstrated its effectiveness. For example, Pitanov et al. (2023) demonstrates its benefits in multi-agent path search, highlighting the advantages over traditional heuristic search methods. Similarly, Li et al. (2023) use MCTS to effectively address various types of SAT problems. Recently, combining MCTS with large-scale language models has been a great trend. Chen et al. (2024) proposed IMCTS, an approach designed to enhance the mathematical reasoning capabilities of fine-tuned LLMs. Xu (2023) integrated MCTS with a lightweight energy function, demonstrating notable performance improvements. In addition, MCTSrDi Zhang et al. (2024) introduced systematic exploration and heuristic self-refinement mechanisms, further advancing its applications in complex decision-making tasks.
Extensive experiments across various downstream applications have consistently validated the efficacy of MCTS in addressing complex planning problems. Building on this success, our work pioneers the application of MCTS in the Text-to-SQL domain for the first time. Our approach incorporates a novel ”fast-and-slow thinking” mechanism, comprising a direct SQL generation module and an MCTS-based refinement module. When direct generation fails, the MCTS module critiques and refines SQL outputs by identifying and addressing errors, iteratively exploring optimal solutions under constrained conditions, thereby reducing hallucination effects caused by the model’s autonomous planning.
3 Preliminaries
This section provides the necessary preliminaries and notations for understanding the proposed framework. We start with a concise overview of the Text-to-SQL task, highlighting its relevance to our approach, and then introduce the fundamental concepts of Monte Carlo Tree Search.
3.1 Definition of Text-to-SQL
Let $\mathcal{X}=(Q,S,\mathcal{K})$ represent the input to the Text-to-SQL task, where $Q$ is a natural language question, $S$ is the database schema, and $\mathcal{K}$ optionally incorporates external knowledge to enhance contextual understanding. The database schema $S$ consists of a set of tables $\mathcal{T}={T_{1},T_{2},\dots,T_{|\mathcal{T}|}}$ and their associated columns $\mathcal{C}={C_{1},C_{2},\dots,C_{|\mathcal{C}|}}$. Each column $C_{i}$ is defined within a specific table $T_{j}$, providing structural and semantic information crucial for SQL generation. The goal of the Text-to-SQL task is to generate an SQL query $\mathcal{Y}$ that precisely captures the semantic intent of the natural language question $Q$ while adhering to the structural constraints of the schema $S$.
3.2 Review of MCTS
Monte Carlo Tree Search is a fundamental tool widely used in game AI and complex planning making. It builds search trees, estimates action values through simulation, and solves complex problems step-by-step through continuous exploration. The algorithm relies on the Upper Confidence bounds for Trees(UCT) to optimize decision efficiency and consists of four key phases: selection, expansion, simulation, and backpropagation.
In the selection phase, the algorithm starts at the root and iteratively selects child nodes based on the UCT until reaching a leaf node. During expansion, if the leaf node is non-terminal, one or more child nodes are added to explore potential future actions. The simulation phase evaluates the newly added nodes by performing stochastic simulations to a terminal state. Finally, in backpropagation, the simulation results results are propagated back through the tree, updating the values of visited nodes and improving future decision-making.
  • Selection: Traverse the tree from the root by selecting child nodes based on UCT until a leaf node is reached.
  • Expansion: If the leaf node is not terminal, extend the tree by adding one or more child nodes to represent unexplored actions.
  • Simulation: Run simulations from the newly added node to a terminal state to estimate its value.
  • Backpropagation: Propagate simulation outcomes back through the visited nodes to update their values and refine the decision policy.
@@ 4 MCTS-SQL Framework
As shown in Figure 2, the MCTS-SQL framework consists of three key components: the Selector, Direct Generator, and MCTS-Refiner. The Selector identifies relevant tables and schema elements based on the user query, while the Direct Generator quickly produces an initial SQL query. Queries that fail or yield errors are refined by the MCTS-Refiner through iterative tree search. A detailed explanation of each component is provided in the subsequent section.
The collaboration process of our MCTS-SQL is presented in Algorithm 1:
@@
8:     while count $<$ maxRollout do
@@
4.1 Schema
Combining the database schema information in the prompt is essential for enabling the LLM to comprehend the database structure accurately and generate precise queries. This paper presents a novel method that illustrates the hierarchical relationships between databases, tables, and columns using a semi-structured format.
To be specific, we provide the table name and corresponding description for each table(which can be omitted if not necessary). The table information is converted into a list where each entry is a tuple containing a column of details. Each column includes the name, data type, description, and example values, thus providing a comprehensive view of its contents. In addition, foreign keys must be included to represent the relationships between tables accurately. Understanding hierarchical relationships is critical for query generation. An example of proposed schema format is shown in Figure 3.
Figure 3: An example of proposed semi-structed database schema format.The format consists of table names, descriptions and column level details(name, data type, description, and examples) to represent the hierarchical information of databeses.
@@ 4.2 Selector
The role of the Selector can be formally described as follows. Given an input triplet $\mathcal{X}=(Q,S,\mathcal{K})$, where $Q$ is the query, $S={T,C}$ is the database schema consisting of tables ($T$) and columns ($C$), and $\mathcal{K}$ denotes the knowledge provided. The Selector aims to identify a minimal subset of tables and columns, denoted as $S^{\prime}={T^{\prime},C^{\prime}}$, which are necessary to answer the query $Q$. The behavior of the Selector is formally defined as follows:
@@
The design of the Selector is motivated by two key considerations: (i) Including unnecessary schema elements in the prompt increases the risk of irrelevant or extraneous items being incorporated into the generated SQL, which can degrade output quality; (ii) Directly utilizing the entire database schema may result in excessively long prompts, which could lead to higher computational costs and potentially exceed the input length limitations of the language model.
4.3 Generator
To balance accuracy and efficiency in Text-to-SQL, we introduce a fast-slow thinking mechanism consisting of a Direct Generator and an MCTS-Refiner. Initially, the user’s input is processed by the Direct Generator, which leverages COT in one prompt to generate SQL queries rapidly. If the executor successfully executes the SQL and the LLM confirms it meets the user’s intent, the result is immediately returned. If an error occurs or the LLM thinks the output is insufficient, the MCTS-Refiner iterative optimization is based on the error feedback and self-criticism. Once reaching the maximum of iterations, the SQL with the highest evaluation value is selected and provided.
4.3.1 A. Direct Generator
The purpose of the Direct Generator is to generate SQL queries directly within a single step through an end-to-end process. It can be described as follows, where $R$ represents the generated SQL query.

### v3

# MCTS-SQL: Light-Weight LLMs can Master the Text-to-SQL through Monte Carlo Tree Search

@@ Abstract
Text-to-SQL is a fundamental yet challenging task in the NLP area, aiming at translating natural language questions into SQL queries. While recent advances in large language models have greatly improved performance, most existing approaches depend on models with tens of billions of parameters or costly APIs, limiting their applicability in resource-constrained environments. For real world, especially on edge devices, it is crucial for Text-to-SQL to ensure cost-effectiveness. Therefore, enabling the light-weight models for Text-to-SQL is of great practical significance. However, smaller LLMs often struggle with complicated user instruction, redundant schema linking or syntax correctness. To address these challenges, we propose MCTS-SQL, a novel framework that uses Monte Carlo Tree Search to guide SQL generation through multi-step refinement. Since the light-weight models’ weak performance of single-shot prediction, we generate better results through several trials with feedback. However, directly applying MCTS-based methods inevitably leads to significant time and computational overhead. Driven by this issue, we propose a token-level prefix-cache mechanism that stores prior information during iterations, effectively improved the execution speed. Experiments results on the SPIDER and BIRD benchmarks demonstrate the effectiveness of our approach. Using a small open-source Qwen2.5-Coder-1.5B, our method outperforms ChatGPT-3.5. When leveraging a more powerful model Gemini 2.5 to explore the performance upper bound, we achieved results competitive with the SOTA. Our findings demonstrate that even small models can be effectively deployed in practical Text-to-SQL systems with the right strategy.
Figure 1: Execution accuracy comparison of MCTS-SQL across some existing methods. The results show that MCTS-SQL significantly enhances the performance of light-weight models, achieving performance comparable to some larger models. And when using Gemini2.5 as the base model, we achieve results competitive with the SOTA.
@@ 1 Introduction
Text-to-SQL is a task aimed at converting natural queries into SQL, which plays a critical role in data analytics and supports a wide range of real-world applications (Survey; 2024survey). Recent advances in LLM (Gemini; GPT4) have significantly improved the performance of Text-to-SQL systems. However, most of these powerful methods (MAC-SQL; MCS; ESQL) rely on extremely huge models or costly APIs, making them expensive and can not be used in resource-constrained environments. For real-world use, especially on edge devices, running efficiently is crucial to keep low resources cost. This raises an important challenge: How can we enable lightweight models to effectively handle Text-to-SQL tasks ?
Most common errors made by these small LLMs is the mistake understanding of users’ intent, wrong schema selection and syntax errors. Due to the poor performance of single-shot prediction, an intuitive way to address these challenges is to conduct a trial-and-feedback mechanism to iteratively optimize the generated SQL. However, feedback without any constraints or guidance is inefficient. Therefore, a more powerful optimization strategy is needed to direct better solutions. Monte Carlo Tree Search (MCTS) has proven to be an efficient tool in decision-making and optimization tasks. Recent studies (Pitanov; li2023; chen2024) have demonstrated that MCTS can be effectively applied to problems requiring iterative improvements. Given its strengths, MCTS presents a practical tool for optimizing SQL generation in Text-to-SQL tasks.
In this paper, we introduce MCTS-SQL, a practical approach that effectively applies MCTS to the Text-to-SQL task. Every components of our design follows one principle: progressively reducing the search space to align with the limited reasoning capacity of light models.
To realize this, we introduce three components:(i)Selector prunes irrelevant schema elements, reducing prompt complexity; (ii)Direct Generator provides a strong initial SQL candidate, avoiding deep search from scratch;(iii)MCTS-Refiner iteratively improves the SQL through guided exploration.
While effective, MCTS-based approaches introduce a new challenge: computation cost. Compared to single-shot prediction, iterative refinement inevitably increases token usage and inference time. Rethinking the whole pipeline of MCTS-SQL, we find that a large amount of inputs of every instructions are highly repetitive (e.g., database schema, field descriptions, few-shot examples). To address the latency from re-computing these inputs, we design a prefix-cache mechanism. Repeated inputs are computed once and cached, during MCTS iterations, only changing feedback and refinements are processed. This approach reduces inference time by 53%.
We evaluate the performance of MCTS-SQL on two widely-used benchmarks: The Spider (SPIDER) and BIRD (BIRD). The results show that MCTS-SQL based on the Qwen-2.5-Coder-Instruct-1.5B (Qwen) outperforms ChatGPT-3.5, and 3B version achieves even better performance than some earlier GPT-4o based methods. Moreover, to explore the boundaries of our algorithm, we evaluate results based on Gemini-2.5. Unsurprisingly, we achieve competitive results, with execution accuracy of 72.91% on BIRD. The comparison across some existing methods can be seen in Figure 1.
@@
  • We propose MCTS-SQL, which apply Monte Carlo Tree Search to Text-to-SQL, demonstrating that small models can effectively handle this complex task.
  • We design a search-space-reducing pipeline with Selector, Direct Generator, and MCTS-Refiner to guide SQL optimization.
  • We introduce a novel prefix-cache mechanism to reduce redundant computation, significantly improving inference efficiency.
  • Experiments on BIRD and SPIDER show that MCTS-SQL enables small models to achieve practical accuracy. For instance, using Qwen-1.5B, our method reaches 40.69% EX on BIRD dev set. When leveraging a stronger model Gemini 2.5, it achieves 72.91% EX on BIRD dev, which is competitive with the current SOTA.
Figure 2: The MCTS-SQL framework consists of three core components: the Selector, the Direct Generator and the MCTS-Refiner. The Selector is used to filter the most relevant tables and columns based on the user’s intent. The Direct Generator aims to produce an initial SQL query. And the MCTS-Refiner is activated when the initial SQL query fails both execution and LLM-based verification checks. which adopts iterative trial-and-feedback optimization to refine the query progressively.
@@ 2.1 Text-to-SQL
Text-to-SQL aims to bridge natural language queries and structured database queries, and numerous approaches are proposed to address its challenges. Early systems, such as LUNAR (LUNAR) and NaLIX (NaLIX), employed rule-based methods that manually crafted grammar rules and heuristics. However, the generalization performance of these methods across different tasks or databases is difficult to guarantee.
The deep learning marked a turning point for Text-to-SQL. End-to-end models like Seq2SQL (Seq2SQL) and SQLNet (SQLNet) directly mapped natural language to SQL but struggled with complex queries, especially those involving nested structures or intricate reasoning. Pre-trained Language Models (PLMs), such as TaBERT (TaBERT) and BERT-SQL (BERT-SQL), enhance cross-domain generalization and improve the accuracy of SQL generation. However, these methods require a certain amount of domain-specific SQL training data, which makes them difficult to land in practical applications
Recently, Large Language Models (LLMs) such as GPT-4 (GPT4), Palm-2 (Palm), and LLaMA (llama) have revolutionized Text-to-SQL tasks. These methods excel in zero-shot and few-shot settings without any extensive training data.(MCS; Chess; ESQL). DAIL-SQL (DAIL-SQL) optimized prompt engineering, focusing on question representation, prompt structure, and example selection to enhance SQL accuracy with minimal supervision. Additionally, MAC-SQL (MAC-SQL) introduced a collaborative framework integrating decomposer, auxiliary selector, and refiner modules for iterative SQL refinement. Recently, Xiyan-SQL(Xiyan), Chase-SQ(chase)L, and DSAIR-SQL(DSAIR) have also achieved improvements in NL2SQL performance through fine-tuning and sophisticated agent engineering. The most similar works, Alpha-SQL(Alpha) and SQL-o1(SQL-o1), perform heuristic, action-level search from scratch, whereas our MCTS-SQL adopts a progressive refinement strategy and leverages a prefix-cache mechanism to minimize computation during iterative exploration.
From the analysis of existing methods, it becomes clear that advancing Text-to-SQL performance relies on the models’ understanding and reasoning capabilities. However, these models are often of large scale and costly, making them impractical for real-world resource-constrained application. Our core motivation is to enable lightweight models to achieve practical performance in Text-to-SQL tasks. To this end, we incorporate Monte Carlo Tree Search (MCTS) to guide the generation.
@@ 2.2 Monte Carlo Tree Search
MCTS is widely used for planning complex problems, and a large number of downstream experiments have demonstrated its effectiveness. For example, (Pitanov) demonstrates its benefits in multi-agent path search, highlighting the advantages over traditional heuristic search methods. Similarly, (li2023) use MCTS to effectively address various types of SAT problems. Recently, combining MCTS with large-scale language models has been a great trend. (chen2024) proposed IMCTS, an approach designed to enhance the mathematical reasoning capabilities of fine-tuned LLMs. (xu2023) integrated MCTS with a lightweight energy function, demonstrating notable performance improvements. In addition, MCTSr (MCTSr) introduced systematic exploration and heuristic self-refinement mechanisms, further advancing its applications in complex decision-making tasks.
Building on these successes, our work is the first to introduce Monte Carlo Tree Search into the Text-to-SQL domain.The core idea is very simple: reduce errors through iterative trial and error. However, naive exhaustive attempts are inefficient and impractical. To address this, we leverage MCTS to find a more efficiently and reliable exploration path. Moreover, we design a prefix-cache to avoid repeated computational overlead.
@@ 3 MCTS-SQL Framework
As shown in Figure 2, the MCTS-SQL framework consists of three key components: the Selector, Direct Generator, and MCTS-Refiner. The Selector filter relevant tables and schema elements based on the user query, while the Direct Generator produces an initial SQL query. Queries that fail or yield errors are refined by the MCTS-Refiner through iterative tree search. A detailed explanation of each component is provided in the subsequent section. The collaboration process of our MCTS-SQL is presented in Algorithm 1.
3.1 Schema
Before introducing the specific components, we would like to describe the special design of effectively translating database structures. Combining the database schema information in the prompt is essential for enabling the LLM to comprehend the database structure accurately and generate precise queries. We present a novel method that illustrates the hierarchical relationships between databases, tables, and columns using a semi-structured format.
To be specific, we provide the table name and corresponding description for each table(which can be omitted if not necessary). The table information is converted into a list where each entry is a tuple containing a column of details. Each column includes the name, data type, description, and example values, thus providing a comprehensive view of its contents. In addition, foreign keys must be included to represent the relationships between tables accurately. Understanding hierarchical relationships is critical for query generation. All the agents in this paper introduce database information through this schema.
@@ 3.1 Schema
8:  while count << maxRollout do
@@ 3.2 Selector
The role of the Selector can be formally described as follows. Given an input triplet $\mathcal{X}=(Q,S,\mathcal{K})$, where $Q$ is the query, S=T,CS={T,C} is the database schema consisting of tables ($T$) and columns ($C$), and $\mathcal{K}$ denotes the knowledge provided. The Selector aims to identify a minimal subset of tables and columns, denoted as $S^{\prime}={T^{\prime},C^{\prime}}$, which are necessary to answer the query $Q$. The behavior of the Selector is formally defined as follows:
@@
The Selector serves as a schema pruning module that leverages a large language model to identify tables and columns relevant to the query. Given $(Q,S,\mathcal{K})$, it interprets the query and associated knowledge to produce a focused subset $S^{\prime}={T^{\prime},C^{\prime}}$ necessary for answering $Q$. This process eliminates irrelevant schema elements, reducing the noise of schema and preventing prompt overflow caused by the full schema. By reducing the context, the Selector improves both the accuracy and efficiency of subsequent SQL generation.
3.3 Direct Generator
The purpose of the Direct Generator is to generate SQL queries directly through an end-to-end process. It can be described as follows, where $R$ represents the generated SQL query.